OpenStudy (anonymous):

A statement Sn about the positive integers is given. Write statements Sk and Sk+1, simplifying Sk+1 completely. Show your work. Sn: 1 ∙ 2 + 2 ∙ 3 + 3 ∙ 4 + . . . + n(n + 1) = [n(n + 1)(n + 2)]/3

OpenStudy (anonymous):

So Sk=k(k+1)(k+2)3 Then we have Sk+1=(k+1)(k+1+1)(k+1+2)3=(k+1)(k+2)(k+3)3 is what i got so far am i right

OpenStudy (anonymous):

and can you help me finish it

OpenStudy (anonymous):

@lυἶცἶ0210

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

@denise19

OpenStudy (anonymous):

@OregonDuck